Assume the random variable x is normally distributed with mean mu equals 50 and standard deviation sigma equals 7. Find the indicated probability. Upper P left parenthesis x greater than 36 right parenthesis Upper P left parenthesis x greater than 36 right parenthesisequals (Round to four decimal places as? needed.)
Answer #1

Solution :

Given that ,

mean = \mu = 50

standard deviation = \sigma = 7

P(X > 36) = 1 - P(x < 36)

= 1 - P((x - \mu ) / \sigma < (36 - 50) / 7)

= 1 - P(z < -2)   Using standard normal table,

= 1 - 0.0228

= 0.9772

P(x > 36) = 0.9772

Probability = 0.9772